How many sums of money, each involving 3 coins, can be formed from a penny, a nickel, a dime, and a quarter?
Answer by Edwin McCravy(20054)

The four coins, a penny, a nickel, a dime, and a quarter is 41 cents.

To get 3 coins you have to take one coin away.

Take away the penny from the four coins and you have 41-1=40 cents
Take away the nickel from the four coins and you have 41-5=36 cents
Take away the dime from the four coins and you have 41-10=31 cents
Take away the quarter from the four coins and you have 41-25=16 cents

That's 4 different sums of money.

Edwin